### v2

# Federated Distributed Key Generation

Abstract
Distributed Key Generation (DKG) is foundational to many threshold-based cryptographic protocols, including threshold signatures, secure multiparty computation, and i-voting systems. However, standard $(t,n)$-DKG protocols typically require a predefined set of $n$ participants and a global threshold $t$, limiting their applicability in open, decentralized settings where network membership is dynamic and node availability can be unpredictable.
We introduce Federated Distributed Key Generation (FDKG), which relaxes these constraints by allowing each participant to select its own ”guardian set” ($k$ nodes, $t$-of-$k$ local threshold) for reconstructing its partial secret key. This mechanism, inspired by Federated Byzantine Agreement, enables flexible trust delegation and enhances resilience. FDKG generalizes traditional DKG and achieves its core functionality in two communication rounds. Our simulation-based security analysis establishes correctness and privacy for generation, and liveness for reconstruction (if an adversary cannot control a participant and kt+1k-t+1 of its guardians). Simulations further confirm FDKG’s efficacy against node unreliability.
In a scenario with 100 potential parties (50 participating, $k=40$ guardians, 80% retention for tallying), FDKG distribution broadcasts total approx. $336$ kB, and reconstruction phase approx. $525$ kB. Both phases scale with $\mathcal{O}(n\cdot k)$ in terms of broadcast data. Groth16 NIZK client-side proving per participant is $\approx 5$ s for distribution, and ranges from $0.6$ s to $29.6$ s for reconstruction, depending on the number of shares processed.
Our work advances distributed cryptography by providing a robust and adaptable DKG primitive suited for dynamic networks, with potential applications ranging from ad-hoc collaborations and decentralized credential management to secure blockchain governance.
I Introduction
Distributed Key Generation (DKG) protocols are essential for establishing shared cryptographic keys in decentralized systems, thereby eliminating single points of trust and failure. They form the bedrock of numerous applications, including threshold signatures, secure multi-party computation (MPC), and privacy-preserving internet voting systems where distributed trust is paramount [1]. However, classical DKG approaches often fall short in highly dynamic or permissionless environments due to their inherent assumptions about network stability and participant knowledge.
A standard $(t,n)$-DKG protocol [2] involves a fixed set of $n$ nodes collectively generating a secret key whose shares are distributed such that any $t$ or more nodes can reconstruct or use the secret, while fewer than $t$ learn nothing. This model typically requires $n$ to be known at the outset to determine polynomial degrees and sharing parameters. In open settings where participation is voluntary (e.g., public blockchains or ad-hoc networks), the actual number of participants may only become clear after an initial interaction phase. If this number deviates significantly from an anticipated $n$, a pre-selected global threshold $t$ might become unachievable [3, 4] or insecure, rendering the system ineffective. Furthermore, traditional DKG often assumes uniform trust among all $n$ participants, which may not hold in diverse decentralized ecosystems.
To overcome these limitations, we introduce Federated Distributed Key Generation (FDKG), a novel protocol that draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols [5], designed for scenarios with uncertain participation and heterogeneous trust relationships. FDKG allows any subset of potential parties to participate in the key generation process. Its core innovation lies in Guardian Sets: each participating party $P_{i}$ independently selects a trusted subset of $k$ other parties (its guardians) and defines a local threshold $t$ necessary for these guardians to reconstruct $P_{i}$’s partial secret key. This federated approach to trust delegation provides enhanced flexibility and resilience.
Formally, an FDKG protocol involves $n$ potential parties. A subset $\mathbb{D}\subseteq\mathbb{P}$ chooses to participate in generating partial secret keys. Each participant $P_{i}\in\mathbb{D}$ distributes its partial secret key $d_{i}$ among its chosen guardian set $\mathcal{G}_{i}$ of size $k$ using a local $(t,k)$-threshold SSS scheme. The global secret key d is the sum of all valid partial secrets $d_{i}$. During reconstruction, d can be recovered if, for each $P_{i}\in\mathbb{D}$, either $P_{i}$ itself provides $d_{i}$, or at least $t$ of its guardians in $\mathcal{G}_{i}$ provide their respective shares. FDKG generalizes traditional $(t,n)$-DKG, which can be viewed as a special case where all $n$ parties participate and each party’s guardian set comprises all other n1n-1 parties.
This novel approach offers several advantages:
  • Dynamic Participation: FDKG accommodates an arbitrary number of participants $P_{i}\in\mathbb{D}$ joining the key generation process without prior global coordination on the total number of participants.
  • Flexible Trust Model: By allowing each participant to define its own guardian set $\mathcal{G}_{i}$ and local threshold $t$, FDKG supports diverse and evolving trust relationships, making it suitable for decentralized settings where trust is not uniform. This is illustrated in Figure 1, which contrasts FDKG’s federated trust with centralized and traditional distributed trust models.
  • Resilience to Node Unavailability: The use of local thresholding via guardian sets ensures that a participant’s contribution to the global key can be recovered even if the participant itself becomes unavailable, provided a sufficient number of its designated guardians remain honest and active.
Figure 1: Comparison of trust models in distributed systems: centralized trust (one authority), distributed trust (multiple authorities), and federated trust (peer-to-peer with individually chosen trusted groups).
In this paper, we provide a detailed specification of the FDKG protocol, a formal security analysis within the simulation-based (Real/Ideal) paradigm (Section V), and demonstrate its practical utility through an application to an internet voting scheme. Our primary contribution is the FDKG framework itself, which enables robust and verifiable key generation with adaptable trust for dynamic networks.
This paper makes the following contributions:
  • We formally define the FDKG protocol, detailing its key distribution and reconstruction phases.
  • We provide a security analysis of FDKG in the simulation-based framework, proving its correctness and privacy for key generation, and analyzing the liveness conditions for key reconstruction (Section V).
  • We demonstrate FDKG’s performance and resilience through extensive simulations under various network conditions, participation rates, and retention scenarios (Section VI), and evaluate the computational and communication costs of its cryptographic components (Section VIII).
  • We integrate FDKG into an end-to-end internet voting scheme, showcasing its practical application using threshold ElGamal encryption and NIZK-based verifiability (Section VII). A corresponding open-source implementation is provided.
The remainder of this paper is organized as follows. Section II reviews related work in DKG. Section III presents the cryptographic primitives. Section IV describes the FDKG protocol. Section V provides the security analysis. Section VI discusses simulation results. Section VII details the voting application. Section VIII assesses performance. Section IX outlines deployment strategies. Finally, Section X discusses limitations and future work, and Section XI offers concluding remarks.
II Related Work
Distributed Key Generation (DKG) is a cornerstone of threshold cryptography, enabling a set of parties to jointly compute a shared public key and individual secret key shares without any single party learning the entire secret key. Seminal DKG protocols, such as those by Gennaro et al. [2], established foundational techniques based on verifiable secret sharing (VSS) [6] over discrete logarithm-based cryptosystems, typically using Shamir’s Secret Sharing (SSS) [7] for share distribution. These protocols ensure that any authorized threshold of parties can reconstruct or use the secret, while unauthorized subsets gain no information.
Subsequent research has extensively explored DKG, focusing on various aspects such as robustness against adversarial behavior, efficiency, round complexity, and applicability in different network models (synchronous vs. asynchronous). Considerable effort has been dedicated to enhancing DKG for reliable operation in challenging environments like the internet, with a focus on asynchronous settings and high thresholds of corruptions [8, 9, 10, 11]. These works primarily address fault tolerance and adversarial resilience within a framework where the set of participants is often known or becomes known during an initial setup phase.
A significant body of work focuses on Publicly Verifiable Secret Sharing (PVSS) [12, 13], where the correctness of shared information can be verified by any observer, not just the share recipients. PVSS is crucial for non-interactive DKG protocols, where parties broadcast their contributions along with proofs, minimizing interaction rounds. Groth’s work on DKG and key resharing [14] presents efficient constructions leveraging pairing-based cryptography and NIZK proofs, aiming for minimal rounds and public verifiability, which is a common goal for modern DKG deployed on transparent ledgers or bulletin boards. The security definitions and expected rigor for DKG protocols, particularly those aiming for provable security in simulation-based models, are exemplified in recent literature such as Katz’s work on round-optimal, fully secure DKG [15].
Our work on Federated Distributed Key Generation (FDKG) addresses a distinct set of challenges primarily related to dynamic participation and flexible trust delegation in settings where the total number of participants is not known beforehand, and individual participants may have heterogeneous trust in others. Traditional DKG and many NI-DKG schemes [2, 14, 15] typically assume a fixed set of $n$ participants for whom the DKG is run, or require mechanisms to establish this set. The polynomial degree in SSS, for example, depends on the intended threshold, which itself is often related to $n$. If $n$ is unknown until after the first round of messages (where parties declare participation), pre-determining a globally optimal polynomial degree becomes problematic.
Dynamic DKG schemes [16] were proposed to allow parties to join and leave an existing DKG setup. While these offer flexibility, they often involve multiple rounds of interaction for updates, which can be cumbersome for applications requiring rapid setup or involving human-in-the-loop interaction, such as ad-hoc collaborations or internet voting applications. FDKG aims to achieve its key generation and initial share distribution within a minimal number of rounds (one for distribution and one for reconstruction).
FDKG introduces participant-defined ”guardian sets,” allowing each key contributor to specify a trusted subset of parties responsible for helping reconstruct their specific partial secret key. This localized trust model, where each participant $P_{i}$ defines its own $(t,k)$-threshold for its share $d_{i}$ among $k$ guardians, contrasts with the global $(t,n)$-threshold common in standard DKG. This approach offers enhanced adaptability in unreliable or permissionless networks where global trust assumptions are hard to establish or maintain. The flexibility offered by FDKG generalizes traditional DKG protocols; a standard $(t,n)$-DKG can be seen as a special case of FDKG where every participant considers all other n1n-1 parties as their guardians.
Our design draws conceptual inspiration from Federated Byzantine Agreement (FBA) protocols, such as the Stellar Consensus Protocol [5]. FBA allows open participation by letting nodes choose their own ”quorum slices” (sets of trusted peers), enabling decentralized consensus without a globally predefined set of validators. Similarly, FDKG allows participants to define their local trust spheres (guardian sets) for key management. As a party is included in more guardian sets, its importance for liveness within the FDKG network naturally increases, reflecting a form of emergent trust or criticality.
The primary motivation for FDKG in this paper is its application to internet voting. Existing e-voting systems often rely on a pre-selected, fixed set of trusted authorities (Key Holders or Trustees) to manage decryption keys [17, 18, 19, 20, 21, 22]. While these systems use threshold cryptography to protect privacy against a subset of colluding authorities, their reliance on a static, globally trusted set makes them vulnerable if these authorities are compromised, collude beyond the threshold, or are simply unavailable during critical phases like tallying. FDKG aims to decentralize this trust by allowing individual participants (or groups acting on their behalf) to establish partial keys and manage their recovery through self-chosen guardians, enhancing system resilience and flexibility against such failures. This approach aligns with the goal of creating more robust and decentralized foundations for secure multi-party applications.
III Preliminaries
This section introduces the cryptographic primitives, concepts, and notations foundational to our Federated Distributed Key Generation (FDKG) protocol.
III-A Notation
Let $\lambda$ denote the computational security parameter. We work within a cyclic group $\mathbb{G}$ of prime order $q$ generated by $G\in\mathbb{G}$. Group operations are written multiplicatively, and scalar multiplication is denoted by $G^{a}$ for $a\in\mathbb{Z}_{q}$. We assume the Discrete Logarithm Problem (DLP) is hard in $\mathbb{G}$. All algorithms are assumed to be Probabilistic Polynomial-Time (PPT) unless stated otherwise. Sampling a random element $x$ from a set $S$ is denoted by $x\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,S$. We denote the set of potential parties as $\mathbb{P}=\{P_{1},\dots,P_{n}\}$.
III-B Shamir’s Secret Sharing (SSS)
A $(t,n)$-threshold Shamir Secret Sharing scheme over the finite field $\mathbb{Z}_{q}$ (where $q$ is prime and $q>n$) allows a dealer to share a secret $s\in\mathbb{Z}_{q}$ among $n$ parties $P_{1},\dots,P_{n}$ such that any subset of $t$ or more parties can reconstruct $s$, while any subset of fewer than $t$ parties learns no information about $s$.
  • $(s_{1},\dots,s_{n})\leftarrow\textrm{Share}(s,t,n)$: The dealer chooses a random polynomial $f(X)=\sum_{k=0}^{t-1}a_{k}X^{k}\in\mathbb{Z}_{q}[X]$ of degree t1t-1 such that $f(0)=s$. For each party $P_{i}$ ($i\in[n]$), the dealer computes the share $s_{i}=f(i)$.
  • $s\leftarrow\textrm{Reconstruct}(I,\{s_{i}\}_{i\in I})$: Given a set $I\subseteq[n]$ of at least $t$ indices and their corresponding shares $\{s_{i}\}_{i\in I}$, this algorithm computes $s=\sum_{i\in I}\lambda_{i}s_{i}$, where $\lambda_{i}$ are the Lagrange coefficients for the set $I$ evaluated at 0.
SSS provides perfect privacy against adversaries controlling fewer than $t$ parties.
III-C Verifiable and Publicly Verifiable Secret Sharing (VSS/PVSS)
VSS schemes enhance SSS by allowing parties to verify the consistency of their shares relative to some public information. PVSS strengthens this by enabling any entity to perform this verification using only public data [12, 13]. PVSS schemes typically involve the dealer encrypting each share $s_{i}$ under the recipient $P_{i}$’s public key $\textrm{pk}_{i}$ and publishing these ciphertexts $C_{i}=\texttt{Enc}(\textrm{pk}_{i},s_{i})$ along with public commitments to the secret polynomial (e.g., $\{A_{k}=G^{a_{k}}\}$) and a public proof $\pi$ of consistency. Our FDKG protocol employs a PVSS scheme where each participant acts as a dealer.
DKG from PVSS: In a typical PVSS-based DKG [15], each party $P_{k}$ acts as a dealer for a partial secret $d_{k}=f_{k}(0)$. $P_{k}$ uses PVSS to distribute shares $d_{k,j}=f_{k}(j)$ to all parties $P_{j}$, publishing commitments $\{E_{k,l}=G^{a_{k,l}}\}$, encrypted shares $\{C_{k,j}\}$, and a proof $\pi_{k}$. The final secret key is $d=\sum_{k}d_{k}$, the public key is $E=\prod_{k}E_{k,0}$, and party $P_{i}$’s share is $d_{i}=\sum_{k}d_{k,i}$.
III-D Public Key Encryption (PKE)
A public-key encryption scheme $\mathcal{E}=(\textrm{KeyGen},\textrm{Enc},\textrm{Dec})$ consists of three algorithms:
  • $(\textrm{pk}_{i},\textrm{sk}_{i})\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{KeyGen}(1^{\lambda})$: Generates a public key $\textrm{pk}_{i}$ and a secret key $\textrm{sk}_{i}$.
  • $c\leftarrow\textrm{Enc}(\textrm{pk}_{i},m,\textrm{rand})$: Encrypts message $m$ under $\textrm{pk}_{i}$ using provided randomness rand to produce ciphertext $c$.
  • $m/\perp\leftarrow\textrm{Dec}(\textrm{sk}_{i},c)$: Decrypts ciphertext $c$ using secret key $\textrm{sk}_{i}$ to recover message $m$ or outputs failure $\perp$.
We require IND-CPA security (Indistinguishable under Chosen Plaintext Attack).
III-E Non-Interactive Zero-Knowledge (NIZK) Proofs
NIZK proofs allow proving the truth of an NP statement $x\in\mathcal{L}$ (for relation $\mathcal{R}$) without interaction and without revealing the witness $w$.
NIZKs require a setup phase. Different models exist regarding the nature and trust assumptions of this setup:
  • Structured Reference String (SRS) / Trusted Setup: Systems like Groth16 [23] require a CRS generated with specific mathematical structure. This structure is essential for proof succinctness and efficiency but typically requires a trusted setup ceremony (or an MPC equivalent) to generate the CRS without revealing a trapdoor. The CRS is specific to a particular circuit/relation.
  • Universal / Updatable Setup: Systems like PLONK [24] or Marlin [25] use a CRS that can be used for any circuit up to a certain size (universal) and may allow the CRS to be updated securely without repeating a full trusted ceremony (updatable). They still require an initial setup phase, often trusted.
  • Transparent Setup / Random Oracle Model (ROM): Systems like STARKs [26] require minimal setup assumptions, often relying only on public randomness (transparent) or operating in the idealized Random Oracle Model (ROM), where a public hash function is modeled as a truly random function. These avoid complex or trusted setup ceremonies.
Our current implementation utilizes Groth16 due to its efficiency and widespread tooling, thus operating in the CRS model and requiring a trusted setup. However, the FDKG framework itself is compatible with other NIZK systems.
A NIZK system (e.g., in the CRS model) provides:
  • $\textrm{crs}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{Setup}(1^{\lambda},\mathcal{R})$: Generates the public CRS. (A simulator might receive additional trapdoor information during security analysis, depending on the scheme).
  • $\pi\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{Prove}(\textrm{crs},x,w)$: Produces proof $\pi$.
  • $b\leftarrow\textrm{Verify}(\textrm{crs},x,\pi)$: Outputs $1$ (accept) or 0 (reject).
We require standard Completeness, computational Soundness, and computational Zero-Knowledge.
III-F Encrypted Channels for Share Distribution
Protocols like PVSS and DKG often require confidential transmission of secret shares (e.g., $s_{i}$) to specific recipients ($P_{i}$) over a public broadcast channel. This is achieved using an IND-CPA secure PKE scheme. Each potential recipient $P_{i}$ has a key pair $(\textrm{pk}_{i},\textrm{sk}_{i})$ where $\textrm{pk}_{i}$ is public. To send $s_{i}$ to $P_{i}$, the sender computes $c=\texttt{Enc}(\textrm{pk}_{i},s_{i},\textrm{rand})$ and broadcasts $c$. Only $P_{i}$ can compute $s_{i}=\texttt{Dec}(\textrm{sk}_{i},c)$.
For circuit efficiency with Groth16, we use an ElGamal variant over BabyJub [27, 28, 29]. The specific ElGamal variant encryption and decryption algorithms used in our implementation are detailed in Algorithm1 and Algorithm2, respectively.
1 Input: Recipient public key $\textrm{pk}_{i}$, scalar message $m\in\mathbb{Z}_{q}$, randomness $(k,r)\in\mathbb{Z}_{q}\times\mathbb{Z}_{q}$
2 Output: Ciphertext tuple $(C_{1},C_{2},\Delta)$
3 $C_{1}=G^{k}$
4 $M=G^{r}$
5 $C_{2}=(\textrm{pk}_{i})^{k}\cdot M$
6 $\Delta=M.x-m\pmod{q}$
return $(C_{1},C_{2},\Delta)$
Algorithm 1 ElGamal Encryption $\texttt{Enc}(\textrm{pk}_{i},m,[k,r])$
1 Input: Recipient secret key $\textrm{sk}_{i}$, ciphertext $(C_{1},C_{2},\Delta)$
2 Output: Scalar message $m\in\mathbb{Z}_{q}$
3 $M=C_{2}\cdot(C_{1}^{\textrm{sk}_{i}})^{-1}$
4 $m=M.x-\Delta\pmod{q}$
return $m$
Algorithm 2 ElGamal Decryption $\texttt{Dec}(\textrm{sk}_{i},(C_{1},C_{2},\Delta))$
III-G Summary of Key Notations
Table I: Key Notations Used in the FDKG Protocol
Notation Description
$n$ Total number of potential parties in the system $\mathbb{P}$.
$\mathbb{P}$ Set of all potential parties $\{P_{1},\dots,P_{n}\}$.
$P_{i}$ Identifier for party $i$.
$\mathbb{G},G,q$ Cyclic group, its generator, and prime order.
$\textrm{pk}_{i},\textrm{sk}_{i}$ PKE public/secret key pair for $P_{i}$ (for encrypted channels).
crs Common Reference String for NIZK proofs.
$\pi$ Generic symbol for a NIZK proof (e.g., $\pi_{FDKG_{i}}$).
$t$ Local threshold for SSS reconstruction within a guardian set.
$k$ Size of each participant’s chosen guardian set.
$\mathbb{D}$ Set of parties participating and validated in the FDKG phase.
$d_{i},E_{i}$ Partial secret/public key of participant $P_{i}\in\mathbb{D}$.
($E_{i}=G^{d_{i}}$).
$\textbf{d},\textbf{E}$ Global secret/public key aggregated from partial keys.
($\textbf{d}=\sum d_{i}$, $\textbf{E}=\prod E_{i}=G^{\textbf{d}}$).
$\mathcal{G}_{i}$ Guardian set chosen by participant $P_{i}$.
$s_{i,j}$ SSS share of $d_{i}$ for guardian $P_{j}\in\mathcal{G}_{i}$.
$C_{i,j}$ Ciphertext of $s_{i,j}$ encrypted under $\textrm{pk}_{j}$.
$\texttt{Enc}(\textrm{pk},\cdot,\cdot)$ PKE encryption algorithm.
$\texttt{Dec}(\textrm{sk},\cdot)$ PKE decryption algorithm.
$\mathbb{V}$ Set of parties participating in the voting phase.
$B_{i},v_{i},r_{i}$ Voter $P_{i}$’s encrypted ballot, encoded vote, and blinding factor.
$\mathbb{T}$ Set of parties participating in the online tally phase.
$C1,C2$ Aggregated ElGamal components of all cast ballots.
$\mathrm{PD}_{i}$ Partial decryption of $C1$ by $P_{i}$ using $d_{i}$.
$\mathrm{PD}_{j}]_{i}$ Partial decryption of $C1$ by guardian $P_{i}$ using share $s_{j,i}$.
IV Federated Distributed Key Generation (FDKG)
Federated Distributed Key Generation (FDKG) extends traditional Distributed Key Generation (DKG) protocols to accommodate settings with dynamic participation and participant-defined trust assumptions. Standard $(t,n)$-DKG protocols require a fixed set of $n$ participants and a global threshold $t$ for secret reconstruction [2]. FDKG relaxes these requirements, allowing any subset of potential participants to engage in key generation, each delegating reconstruction authority to a self-chosen ”guardian set”.
IV-A Protocol Setting and Goal
We consider a set $\mathbb{P}=\{P_{1},\dots,P_{n}\}$ of $n$ potential parties. The system operates over a cyclic group $\mathbb{G}$ of prime order $q$ with generator $G$. We assume a Public Key Infrastructure (PKI) where each party $P_{i}$ possesses a long-term secret key $\textrm{sk}_{i}$ and a corresponding public key $\textrm{pk}_{i}$ for an IND-CPA secure public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ (Section III-D), used for secure share distribution. Communication occurs via a broadcast channel. The protocol also utilizes a Non-Interactive Zero-Knowledge (NIZK) proof system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ (Section III-E) which is assumed to be complete, computationally sound, and zero-knowledge, operating with a common reference string crs.
The goal of FDKG is for any subset of participating parties $\mathbb{D}\subseteq\mathbb{P}$ to jointly generate a global public key $\textbf{E}\in\mathbb{G}$ such that the corresponding secret key $\textbf{d}\in\mathbb{Z}_{q}$ (where $\textbf{E}=G^{\textbf{d}}$) is implicitly shared among them. Specifically, each participant $P_{i}\in\mathbb{D}$ generates a partial key pair $(d_{i},E_{i})$ where $E_{i}=G^{d_{i}}$. The global key pair is formed by aggregation: $\textbf{d}=\sum_{i\in\mathbb{D}}d_{i}$ and $\textbf{E}=\prod_{i\in\mathbb{D}}E_{i}$. Each participant $P_{i}$ uses a local $(t,k)$-Shamir Secret Sharing (SSS) scheme (Section III-B) to distribute its partial secret $d_{i}$ among a self-selected guardian set $\mathcal{G}_{i}\subset\mathbb{P}\setminus\{P_{i}\}$ of size $k$.
IV-B The FDKG Protocol ($\Pi_{FDKG}$)
The protocol proceeds in two rounds: Distribution and Reconstruction.
IV-B1 Round 1: Distribution Phase
Any party $P_{i}\in\mathbb{P}$ may (optionally) participate:
  1. 1.
    Guardian Selection: Choose $\mathcal{G}_{i}\subset\mathbb{P}\setminus\{P_{i}\}$ with $|\mathcal{G}_{i}|=k$.
  2. 2.
    Partial Key Generation: Sample partial secret $d_{i}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}$. Compute $E_{i}=G^{d_{i}}$.
  3. 3.
    Share Generation: Generate shares $(s_{i,1},\dots,s_{i,n})\leftarrow\textrm{Share}(d_{i},t,n)$.
  4. 4.
    Share Encryption: Initialize empty lists for ciphertexts $\mathbb{C}_{i}$ and randomness $\mathbb{R}_{i}$. For each $P_{j}\in\mathcal{G}_{i}$:
    • Sample PKE randomness: $(k_{i,j},r_{i,j})\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}\times\mathbb{Z}_{q}$.
    • Encrypt: $C_{i,j}\leftarrow\texttt{Enc}_{\textrm{pk}_{j}}(s_{i,j},[k_{i,j},r_{i,j}])$.
    • Store: Add $C_{i,j}$ to $\mathbb{C}_{i}$ and add $(k_{i,j},r_{i,j})$ to $\mathbb{R}_{i}$.
  5. 5.
    Proof Generation: Let the statement be $x_{i}=(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{C}_{i})$. Let the witness be $w_{i}=(d_{i},\{s_{i,j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{R}_{i})$. Generate $\pi_{FDKG_{i}}\leftarrow\textrm{Prove}(\textrm{crs},x_{i},w_{i})$.
  6. 6.
    Broadcast: Broadcast $(P_{i},E_{i},\mathcal{G}_{i},\mathbb{C}_{i},\pi_{FDKG_{i}})$.
Post-Round 1 Processing (by any observer):
  • Initialize $\mathbb{D}=\emptyset$.
  • For each received tuple $(P_{i},E_{i},G_{i},\mathbb{C}_{i},\pi_{i})$:
    • Construct statement $x_{i}=(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in G_{i}},\mathbb{C}_{i})$.
    • Verify proof $b\leftarrow\textrm{Verify}(\textrm{crs},x_{i},\pi_{i})$.
    • If $b=1$, add $P_{i}$ to $\mathbb{D}$ and store $(E_{i},\mathcal{G}_{i},\mathbb{C}_{i})\leftarrow(E_{i},G_{i},\mathbb{C}_{i})$.
  • Compute $\textbf{E}=\prod_{P_{i}\in\mathbb{D}}E_{i}$.
  • Public state: $\mathbb{D}$, E, $\{E_{i},\mathcal{G}_{i},\mathbb{C}_{i}\}_{P_{i}\in\mathbb{D}}$.
IV-B2 Round 2: Reconstruction Phase
Online Reconstruction (by parties $P_{i}\in\mathbb{T}\subseteq\mathbb{P}$):
  1. 1.
    Reveal Partial Secret: If $P_{i}\in\mathbb{D}\cap\mathbb{T}$:
    • Generate $\pi_{PS_{i}}\leftarrow\textrm{Prove}(\textrm{crs},E_{i},d_{i})$.
    • Broadcast $(P_{i},\text{secret},d_{i},\pi_{PS_{i}})$.
  2. 2.
    Reveal Received Shares: For each $C_{j,i}$ (where $P_{j}\in\mathbb{D},P_{i}\in\mathcal{G}_{j}$):
    • Decrypt share: $s_{j,i}\leftarrow\texttt{Dec}(\textrm{sk}_{i},C_{j,i})$.
    • Generate proof: $\pi_{SPS_{j,i}}\leftarrow\textrm{Prove}(\textrm{crs},(C_{j,i},\textrm{pk}_{i},s_{j,i}),\textrm{sk}_{i})$.
    • Broadcast $(P_{i},\text{share},P_{j},s_{j,i},\pi_{SPS_{j,i}})$.
Offline Reconstruction (by any observer):
  1. 1.
    Initialize empty list $D=\emptyset$.
  2. 2.
    For each participant $P_{i}\in\mathbb{D}$:
    • Direct Secrets Reconstruction: Look for $(P_{i},\text{secret},d_{i},\pi_{PS_{i}})$. If $1=\textrm{Verify}(\textrm{crs},E_{i},\pi_{PS_{i}})$, add $d_{i}$ to $D$;
    • Share-Based Reconstruction
      1. (a)
        Initialize empty list $S_{i}$. For each $(P_{k},\text{share},P_{i},s_{i,j},\pi_{SPS_{i,j}})$ where $P_{k}=P_{j}$ and $C_{i,j}$ exists: Verify $b\leftarrow\textrm{Verify}(\textrm{crs},(C_{i,j},\textrm{pk}_{j},s_{i,j}),\pi_{SPS_{i,j}})$. If $b=1$, add $(P_{j},s_{i,j})$ to $S_{i}$.
      2. (b)
        Let $I_{i}=\{P_{j}\mid(P_{j},\cdot)\in S_{i}\}$. If $|I_{i}|\geq t$: Select $I^{\prime}_{i}\subseteq I_{i}$, $|I^{\prime}_{i}|=t$. Compute $d_{i}\leftarrow\textrm{Reconstruct}(I^{\prime}_{i},\{s_{i,j}\mid(P_{j},s_{i,j})\in S_{i}\text{ and }P_{j}\in I^{\prime}_{i}\})$. If $G^{d_{i}}=E_{i}$, add $d_{i}$ to $D$.
  3. 3.
    Compute Global Key: If $D$ contains a valid entry for every $P_{i}\in\mathbb{D}$, compute $\textbf{d}=\sum_{d_{i}\in D}d_{i}\pmod{q}$. Output d. Else output failure.
IV-C Communication Complexity
The communication complexity depends on the number of participants $|\mathbb{D}|$ and the size of guardian sets $k$.
  • Round 1 (Distribution): Each participant $P_{i}\in\mathbb{D}$ broadcasts its partial public key $E_{i}$ (1 group element), $k$ encrypted shares $\{C_{i,j}\}$ (size depends on PKE scheme, in our scheme $k\times(2|\mathbb{G}|+|\mathbb{Z}_{q}|)$), and one NIZK proof $\pi_{FDKG_{i}}$. Assuming $|\mathbb{D}|\approx n$, the complexity is roughly $O(nk\cdot|C_{i,j}|+n\cdot|\pi_{FDKG}|)$. In our BabyJub/Groth16 instantiation, this leads to $O(nk)$ complexity dominated by encrypted shares.
  • Round 2 (Online Reconstruction): Each online party $P_{i}\in\mathbb{T}$ might broadcast its partial secret $d_{i}$ (1 scalar) and proof $\pi_{PS_{i}}$, plus potentially many decrypted shares $s_{j,i}$ (1 scalar each) and proofs $\pi_{SPS_{j,i}}$. Let $m_{i}$ be the number of shares party $P_{i}$ reveals (as a guardian for others). The complexity is roughly $O(|\mathbb{T}|\cdot(|\mathbb{Z}_{q}|+|\pi_{PS_{i}}|+m_{max}\cdot(|\mathbb{Z}_{q}|+|\pi_{SPS_{j,i}}|)))$, where $m_{max}=\max_{i}m_{i}$. In the worst case, where $|\mathbb{T}|\approx n$ and $m_{max}\approx n$, this can approach $O(n^{2})$.
FDKG offers $O(n\cdot k)$ complexity for distribution, which is similar to standard DKG if $k\approx n$. Reconstruction complexity varies but can reach $O(n^{2})$ in worst-case guardian configurations.
IV-D FDKG Example
Consider $n=10$ potential parties $\{P_{1},\dots,P_{10}\}$, local threshold $t=2$, and guardian set size $k=3$. Suppose $\mathbb{D}=\{P_{1},P_{3},P_{5},P_{7},P_{9}\}$ participate validly in Round 1.
Figure 2: Visual comparison of $(k=10,t=5)$-DKG protocol (above) and $(n=10,k=3,t=2)$-FDKG (below)
  • Round 1 (as per Section IV-B1): $P_{1}$ chooses $\mathcal{G}_{1}=\{P_{2},P_{3},P_{5}\}$, generates $(d_{1},E_{1})$, shares $d_{1}$ into $s_{1,2},s_{1,3},s_{1,5}$, encrypts them to get $\mathbb{C}_{1}=\{C_{1,2},C_{1,3},C_{1,5}\}$, generates $\pi_{FDKG_{1}}$, and broadcasts. Other participants in $\mathbb{D}$ do similarly. Observers compute $\textbf{E}=\prod_{P_{i}\in\mathbb{D}}E_{i}$.
  • Round 2 (Reconstruction, as per Section IV-B2): Suppose $\mathbb{T}=\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{7}\}$.
    • $P_{1}$ broadcasts $(P_{1},\text{secret},d_{1},\pi_{PS_{1}})$.
    • $P_{2}$ (guardian for $P_{1}$) decrypts $C_{1,2}$ (using $\textrm{sk}_{2}$) to get $s_{1,2}$. Broadcasts $(P_{2},\text{share},P_{1},s_{1,2},\pi_{SPS_{1,2}})$.
    • (Similar actions for other talliers and shares).
  • Offline Reconstruction: An observer collects all valid broadcasted secrets and shares.
    • Recovers $d_{1}$ directly.
    • For $P_{5}$ (assuming $P_{5}\notin\mathbb{T}$ or didn’t reveal $d_{5}$): If $P_{3}$ (in $\mathcal{G}_{5}$) revealed $s_{5,3}$ and $P_{7}$ (in $\mathcal{G}_{5}$) revealed $s_{5,7}$ (both with valid proofs), the observer uses these two shares ($t=2$) to $\textrm{Reconstruct}(\{P_{3},P_{7}\},\{s_{5,3},s_{5,7}\})$ yielding $d_{5}^{\prime}$. Verifies $G^{d_{5}^{\prime}}=E_{5}$.
    • After recovering all $d_{i}$ for $P_{i}\in\mathbb{D}$, compute $\textbf{d}=\sum d_{i}$.
V Security Analysis
In this section, we provide a formal security analysis of the Federated Distributed Key Generation (FDKG) protocol ($\Pi_{FDKG}$). We adopt the standard simulation-based security framework commonly used for Distributed Key Generation (DKG) and Multi-Party Computation (MPC) protocols [15, 30], addressing the need for rigorous definitions and proofs aligned with cryptographic literature. This framework involves comparing the execution of the real protocol in the presence of an adversary with an ideal execution involving a simulator and an ideal functionality.
V-A Security Model
  • Parties: We consider a set of $n$ potential parties $\mathbb{P}=\{P_{1},\dots,P_{n}\}$.
  • Communication Model: Parties communicate via a secure broadcast channel (emulating a bulletin board where messages are public and authenticated). Communication occurs in synchronous rounds.
  • Adversary: We consider a static Probabilistic Polynomial-Time (PPT) adversary $\mathcal{A}$ who corrupts a subset $\mathcal{C}\subset\mathbb{P}$ of parties at the beginning of the protocol execution. The adversary controls the corrupted parties entirely, receiving their inputs and determining their messages. Let $|\mathcal{C}|$ be the number of corrupted parties.
  • Participation: Participation in the FDKG generation phase is optional. The set of actual participants $\mathbb{D}\subseteq\mathbb{P}$ emerges during the protocol execution.
  • Setup Requirements: The protocol requires a pre-existing Public Key Infrastructure (PKI) where each party $P_{i}$ has published an authenticated public encryption key $\textrm{pk}_{i}$ (with corresponding secret key $\textrm{sk}_{i}$) for the private channel mechanism (Section III-F). Additionally, the use of NIZK (specifically Groth16 in our implementation) presupposes the existence of a Common Reference String (CRS), crs, generated via a trusted setup procedure or an equivalent setup mechanism.
  • Cryptographic Assumptions: The security proofs rely on the standard cryptographic assumptions associated with the underlying primitives:
    • PKE Security: The public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ used for private channels must be IND-CPA secure.
    • NIZK Security: The NIZK system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ must provide computational soundness and zero-knowledge.
    • Implicitly, security relies on the hardness of the Discrete Logarithm Problem (DLP) in the underlying elliptic curve group $\mathbb{G}$.
  • Security Definition: Security is defined using the Real/Ideal paradigm [31, 32]. We define an ideal functionality $\mathcal{F}_{FDKG}$ that captures the desired properties of FDKG (see Figure 3). The real protocol $\Pi_{FDKG}$ securely realizes $\mathcal{F}_{FDKG}$ if for any PPT adversary $\mathcal{A}$ controlling a set $\mathcal{C}$ in the real world, there exists a PPT simulator $\mathcal{S}$ controlling $\mathcal{C}$ in the ideal world such that no PPT environment $\mathcal{Z}$ can distinguish between interacting with $\mathcal{A}$ and the real protocol, or with $\mathcal{S}$ and the ideal functionality $\mathcal{F}_{FDKG}$.
V-B Ideal Functionality $\mathcal{F}_{FDKG}$
The ideal functionality $\mathcal{F}_{FDKG}$ interacts with the parties $P_{1},\dots,P_{n}$ and a simulator $\mathcal{S}$. It is parameterized by system parameters (group $\mathbb{G}$, generator $G$), FDKG parameters $(t,k)$, and dynamically tracks participation. Its formal description is provided in Figure 3.
Ideal Functionality $\mathcal{F}^{n,t,k}_{FDKG}$
  • 1.
    Initialization:
    • Waits to receive (Init, $\mathcal{C}$) from the simulator $\mathcal{S}$. Stores the set of corrupted parties $\mathcal{C}$. Let $\mathcal{H}=\mathbb{P}\setminus\mathcal{C}$.
    • Initializes an empty set of participants $\mathbb{D}=\emptyset$.
    • Initializes empty storage for polynomials $f_{i}$ and guardian sets $\mathcal{G}_{i}$ for all $i\in\mathbb{P}$.
    • Initializes state flag ‘phase‘ = ’activation’.
  • 2.
    Honest Party Activation: Upon receiving (ActivateHonest, $i,\mathcal{G}_{i}$) from the simulator $\mathcal{S}$ (indicating honest party $P_{i}\in\mathcal{H}$ decides to participate):
    • If ‘phase‘ is ’activation’ and $P_{i}\notin\mathbb{D}$:
      • Add $P_{i}$ to $\mathbb{D}$.
      • Store the provided guardian set $\mathcal{G}_{i}$ (verify $|\mathcal{G}_{i}|=k$, $P_{i}\notin\mathcal{G}_{i}$).
      • Choose a random polynomial $f_{i}(X)$ of degree t1t-1 over $\mathbb{Z}_{q}$ and store it.
      • Send (Activated, $i$) back to the simulator $\mathcal{S}$.
  • 3.
    Corrupt Party Activation: Upon receiving (ActivateCorrupt, $i,f_{i},\mathcal{G}_{i}$) from the simulator $\mathcal{S}$ (on behalf of $P_{i}\in\mathcal{C}$):
    • If ‘phase‘ is ’activation’ and $P_{i}\notin\mathbb{D}$:
      • Add $P_{i}$ to $\mathbb{D}$.
      • Verify degree of $f_{i}$ is t1t-1, $|\mathcal{G}_{i}|=k$, $P_{i}\notin\mathcal{G}_{i}$. Store valid $f_{i}$ and $\mathcal{G}_{i}$.
  • 4.
    Key Computation Trigger: Upon receiving (ComputeKeys) from the simulator $\mathcal{S}$ [^1] Note on Round Termination: For simplicity in the ideal functionality $\mathcal{F}_{FDKG}$, we model the transition from the activation phase to the key computation phase via this ‘(ComputeKeys)‘ trigger sent by the simulator. In a real-world deployment, particularly on a blockchain, the end of the distribution phase would typically be determined by external environmental factors like block height or a predetermined time duration, not controlled by the adversary. Participants submitting messages after this cutoff would be excluded. Our security analysis focuses on the correctness and privacy properties based on valid participation within the allowed timeframe, assuming the round termination mechanism itself is robust. The core security guarantees derived from the underlying cryptography (IND-CPA PKE, NIZK soundness/ZK) are not fundamentally affected by this modeling simplification.:
    • If ‘phase‘ is ’activation’:
      • Set ‘phase‘ = ’computed’.
      • If $\mathbb{D}=\emptyset$: Send (KeysComputed, null, $\emptyset$, $\emptyset$, $\emptyset$, $\emptyset$) to $\mathcal{S}$ and stop this trigger processing.
      • For each $P_{k}\in\mathbb{D}$: Define $d_{k}:=f_{k}(0)$ and $E_{k}:=G^{d_{k}}$. Let $\mathbb{E}=\{E_{k}\}_{k\in\mathbb{D}}$.
      • Compute global secret key $\textbf{d}:=\sum_{k\in\mathbb{D}}d_{k}\pmod{q}$ and global public key $\textbf{E}:=G^{\textbf{d}}$.
      • For each $k\in\mathbb{D}$ and each $j$ s.t. $P_{j}\in\mathcal{G}_{k}$: Compute share $d_{k}]_{j}:=f_{k}(j)$.
      • Send (KeysComputed, E, $\mathbb{E}$, $\{\mathcal{G}_{k}\}_{k\in\mathbb{D}\cap\mathcal{H}}$, $\{(d_{k},f_{k},\mathcal{G}_{k})\}_{k\in\mathbb{D}\cap\mathcal{C}}$, $\{[d_{k}]_{j}\mid k\in\mathbb{D},P_{j}\in\mathcal{G}_{k}\cap\mathcal{C}\}$) to the simulator $\mathcal{S}$.
      • For each honest participant $P_{i}\in\mathbb{D}\cap\mathcal{H}$: Send (Output, E, $\mathbb{E}$, $d_{i}$, $\mathcal{G}_{i}$, $\{[d_{k}]_{i}\mid k\in\mathbb{D},P_{i}\in\mathcal{G}_{k}\}$) to $P_{i}$.
Figure 3: Ideal functionality for Federated Distributed Key Generation $\mathcal{F}^{n,t,k}_{FDKG}$.
V-C Security of the FDKG Protocol
We claim that our FDKG protocol ($\Pi_{FDKG}$, described in Section IV) securely realizes the ideal functionality $\mathcal{F}_{FDKG}$.
Theorem 1 (Security of FDKG Generation).
Assume the public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ used for the private channel is IND-CPA secure, and the NIZK $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ is zero-knowledge and computationally sound. Then, the protocol $\Pi_{FDKG}$ securely realizes the ideal functionality $\mathcal{F}_{FDKG}$ against static adversaries corrupting any set $\mathcal{C}\subset\mathbb{P}$.
Proof Sketch (Simulation Argument).
We demonstrate the security of $\Pi_{FDKG}$ by constructing a PPT simulator $\mathcal{S}$ that interacts with the ideal functionality $\mathcal{F}_{FDKG}$ (Figure 3) and any PPT adversary $\mathcal{A}$ attacking the real protocol $\Pi_{FDKG}$. The goal is to show that the view of $\mathcal{A}$ when interacting with $\mathcal{S}$ in the ideal world is computationally indistinguishable from its view when interacting with honest parties running $\Pi_{FDKG}$ in the real world. This implies that $\Pi_{FDKG}$ securely realizes $\mathcal{F}_{FDKG}$.
The simulator $\mathcal{S}$ internally simulates the execution environment for $\mathcal{A}$. It proceeds as follows:
  1. 1.
    Setup: $\mathcal{S}$ receives the corrupted set $\mathcal{C}$ from $\mathcal{A}$ and forwards it to $\mathcal{F}_{FDKG}$ via (Init, $\mathcal{C}$). $\mathcal{S}$ receives PKE public keys $\{\textrm{pk}_{i}\}_{i\in\mathcal{C}}$ from $\mathcal{A}$ and generates honest key pairs $(\textrm{pk}_{i},\textrm{sk}_{i})$ for $i\in\mathcal{H}$. $\mathcal{S}$ provides $\mathcal{A}$ with all public PKE keys $\{\textrm{pk}_{i}\}_{i\in\mathbb{P}}$.
  2. 2.
    Simulating Round 1 (Distribution Phase):
    • Corrupt Party Actions: When $\mathcal{A}$ decides a corrupt party $P_{i}\in\mathcal{C}$ should participate and broadcasts its message $(E_{i},\{C_{i,j}\}_{j\in\mathcal{G}_{i}},\pi_{FDKG_{i}})$:
      • $\mathcal{S}$ verifies the proof $\pi_{FDKG_{i}}$ using the public parameters and statement components derived from the message. If invalid, $\mathcal{S}$ flags this internally.
      • $\mathcal{S}$ extracts the claimed polynomial $f_{i}$ and guardian set $\mathcal{G}_{i}$ from $\mathcal{A}$’s internal state or message.
      • $\mathcal{S}$ sends (ActivateCorrupt, $i,f_{i},\mathcal{G}_{i}$) to $\mathcal{F}_{FDKG}$.
    • Simulating Honest Party Actions: $\mathcal{S}$ must decide which honest parties $P_{i}\in\mathcal{H}$ participate in this round. This decision can be based on instructions from the environment or a default strategy (e.g., all honest parties participate). For each honest party $P_{i}$ chosen to participate:
      • $\mathcal{S}$ selects a guardian set $\mathcal{G}_{i}$ for $P_{i}$ (consistent with honest behavior, e.g., randomly chosen from $\mathbb{P}\setminus\{P_{i}\}$).
      • $\mathcal{S}$ informs the ideal functionality by sending (Activate, $\mathcal{G}_{i}$) on behalf of $P_{i}$. $\mathcal{S}$ subsequently receives confirmation (Activated, $i$) from $\mathcal{F}_{FDKG}$.
    • Triggering Key Computation in Ideal World: Once $\mathcal{S}$ determines that all parties wishing to participate (from $\mathcal{A}$’s perspective) have declared their intent (i.e., corrupt parties have sent messages, honest parties’ participation decided by $\mathcal{S}$), $\mathcal{S}$ sends (ComputeKeys) to $\mathcal{F}_{FDKG}$.
    • Receiving Ideal Outputs: $\mathcal{S}$ receives the message (KeysComputed, E, $\{E_{k}\}_{k\in\mathbb{D}}$, $\{\mathcal{G}_{k}\}_{k\in\mathbb{D}\cap\mathcal{H}}$, corrupt party details, $\{[d_{k}]_{j}\mid k\in\mathbb{D},P_{j}\in\mathcal{G}_{k}\cap\mathcal{C}\}$) from $\mathcal{F}_{FDKG}$. This defines the set $\mathbb{D}$ of valid participants whose contributions were accepted by the ideal functionality.
    • Constructing and Delivering Simulated Broadcasts: For each honest party $P_{i}\in\mathbb{D}\cap\mathcal{H}$, $\mathcal{S}$ constructs a simulated broadcast message $(P_{i},E_{i},\mathcal{G}_{i},\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}},\pi_{FDKG_{i}}^{sim})$ where:
      • $E_{i}$ and $\mathcal{G}_{i}$ are taken from the information received from $\mathcal{F}_{FDKG}$.
      • $\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}}$ consists of ciphertexts encrypting dummy values (e.g., 0). For each $P_{j}\in\mathcal{G}_{i}$, $\mathcal{S}$ samples PKE randomness $\textrm{rand}_{i,j}$ and computes $C_{i,j}\leftarrow\texttt{Enc}(\textrm{pk}_{j},0,\textrm{rand}_{i,j})$.
      • $\pi_{FDKG_{i}}^{sim}$ is a simulated NIZK proof generated using the NIZK simulator: $\pi_{FDKG_{i}}^{sim}\leftarrow\textrm{Sim}_{\textrm{NIZK}}(\textrm{crs},(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}},\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}}))$.
      $\mathcal{S}$ then presents the complete set of broadcast messages (real ones from $\mathcal{A}$ for corrupt $i\in\mathbb{D}\cap\mathcal{C}$, and simulated ones $M_{i}^{sim}$ for honest $i\in\mathbb{D}\cap\mathcal{H}$) to $\mathcal{A}$.
  3. 3.
    Final Output: $\mathcal{S}$ runs $\mathcal{A}$ until it terminates and outputs whatever $\mathcal{A}$ outputs.
Indistinguishability Argument: The core of the security proof is arguing that the view provided to $\mathcal{A}$ by $\mathcal{S}$ in the ideal world is computationally indistinguishable from $\mathcal{A}$’s view in the real world. This relies on the security properties of the underlying cryptographic primitives:
  • Simulated vs. Real NIZK Proofs: The proofs $\pi_{FDKG_{i}}$ generated by honest parties in the real world are replaced by simulated proofs $\pi_{FDKG_{i}}^{sim}$ generated by $\mathcal{S}$ using the NIZK simulator. The zero-knowledge property of the NIZK system guarantees that $\mathcal{A}$ cannot distinguish between real and simulated proofs.
  • Simulated vs. Real Encrypted Shares: Consider the ciphertexts $\{C_{i,j}\}$ broadcast by an honest party $P_{i}$.
    • If the recipient $P_{j}$ is also honest ($P_{j}\in\mathcal{H}$), $\mathcal{A}$ does not possess the secret key $\textrm{sk}_{j}$ required for decryption. In the real world, $C_{i,j}$ encrypts the real share $s_{i,j}$. In the simulation, it encrypts a dummy value (e.g., 0). By the IND-CPA security of the PKE scheme, $\mathcal{A}$ cannot distinguish these ciphertexts.
    • If the recipient $P_{j}$ is corrupt ($P_{j}\in\mathcal{C}$), $\mathcal{A}$ possesses $\textrm{sk}_{j}$. In the real world, $C_{i,j}$ encrypts the real share $s_{i,j}$. In the simulation, $\mathcal{F}_{FDKG}$ provides the actual share $s_{i,j}$ to $\mathcal{S}$. $\mathcal{S}$ can use this $s_{i,j}$ to compute the ”correct” ciphertext $C_{i,j}$ using fresh randomness and provide this to $\mathcal{A}$ in the simulated broadcast message $M_{i}^{sim}$. Alternatively, if $\mathcal{S}$ encrypts a dummy value, the indistinguishability relies on the combination of IND-CPA and the fact that the overall view remains consistent due to the later steps. The crucial aspect is that $\mathcal{S}$ can provide $\mathcal{A}$ with exactly the information it is entitled to (shares sent to corrupted parties).
  • Consistency of Public Keys: The partial public keys $E_{i}$ for honest participants are provided by $\mathcal{F}_{FDKG}$ based on its internal random polynomial choices. For corrupt participants, $\mathcal{A}$ provides $E_{i}$ along with $\pi_{FDKG_{i}}$. The computational soundness of the NIZK ensures that $\mathcal{A}$ cannot produce a valid proof $\pi_{FDKG_{i}}$ for a pair $(E_{i},\mathbb{C}_{i})$ that is inconsistent with some underlying polynomial $f_{i}(X)$ (up to negligible probability). Therefore, the set of partial keys $\{E_{i}\}_{i\in\mathbb{D}}$ and the derived global key E have computationally indistinguishable distributions in the real and ideal worlds.
  • Participation and Guardian Sets: The process by which the set of valid participants $\mathbb{D}$ is determined (based on valid proofs) and how guardian sets $\mathcal{G}_{i}$ for honest parties are chosen and communicated is consistent between the simulation and the real protocol execution flow.
Combining these points, we conclude that any PPT adversary $\mathcal{A}$’s view in the simulation is computationally indistinguishable from its view in the real protocol execution, assuming the underlying PKE scheme is IND-CPA secure and the NIZK system is zero-knowledge and computationally sound. This completes the proof sketch. ∎
V-D FDKG Security Properties
The simulation-based security theorem (Theorem 1) implies the following essential security properties for the FDKG generation phase (Round 1):
  • Correctness: For the set of valid participants $\mathbb{D}$, the protocol outputs are correctly formed according to the protocol specification. That is, the computed global public key E corresponds to the sum of the implicit partial secret keys $\textbf{E}=G^{\sum_{i\in\mathbb{D}}d_{i}}$, each partial public key relates correctly to its partial secret key $E_{i}=G^{d_{i}}$, and the encrypted shares $C_{i,j}$ correspond to the correct evaluations $f_{i}(j)$ of the underlying polynomial $f_{i}(X)$ defining $d_{i}=f_{i}(0)$. This is guaranteed by the definition of $\mathcal{F}_{FDKG}$ in the ideal world and by the computational soundness of the NIZK proofs $\pi_{FDKG_{i}}$ in the real world.
  • Privacy: The adversary $\mathcal{A}$ (controlling the set $\mathcal{C}$) learns no information about the partial secret keys $d_{i}$ of honest participants ($i\in\mathcal{H}\cap\mathbb{D}$) or the final global secret key d, beyond what can be inferred from the public information (E, $\{E_{i}\}_{i\in\mathbb{D}}$) and the shares explicitly sent to corrupted parties ($\{[d_{i}]_{j}\mid P_{j}\in\mathcal{G}_{i}\cap\mathcal{C}\}$). This follows from the indistinguishability argument in the proof sketch: the simulator $\mathcal{S}$ constructs $\mathcal{A}$’s view without access to honest parties’ secrets, relying on the IND-CPA security of the PKE and the zero-knowledge property of the NIZK system.
  • Robustness (During Generation): Honest participating parties ($\mathcal{H}\cap\mathbb{D}$) successfully complete the key generation phase and obtain their outputs (their partial secret $d_{i}$, the global public key E, etc.) as defined by the ideal functionality $\mathcal{F}_{FDKG}$, irrespective of the (mis)behavior of corrupted parties. Malicious parties $P_{j}\in\mathcal{C}$ cannot prevent protocol completion for honest parties other than by failing to participate or by broadcasting messages with invalid proofs $\pi_{FDKG_{j}}$, which are simply ignored by observers (as reflected in the Post-Round 1 processing).
V-E Liveness of Key Reconstruction
Distinct from the security properties of the generation phase, Liveness addresses the success of the reconstruction phase (Round 2). It guarantees that the global secret key d can be recovered if enough participants and their guardians are available and behave honestly during this second phase. This property assumes the generation phase successfully realized $\mathcal{F}_{FDKG}$.
Definition 1 (Liveness of Reconstruction).
The FDKG protocol achieves Liveness of Reconstruction if, for every participant $P_{i}\in\mathbb{D}$ whose contribution $d_{i}$ is required for an application using the key, at least one of the following conditions holds during the reconstruction phase:
  • $P_{i}$ is honest and available, and successfully broadcasts $(P_{i},\text{secret},d_{i},\pi_{PS_{i}})$ where $\pi_{PS_{i}}$ is valid.
  • At least $t$ distinct parties $\{P_{j_{1}},\dots,P_{j_{t}}\}\subseteq\mathcal{G}_{i}$ are honest and available, and each successfully broadcasts $(P_{j_{r}},\text{share},P_{i},s_{i,j_{r}},\pi_{SPS_{i,j_{r}}})$ where $\pi_{SPS_{i,j_{r}}}$ is valid.
Liveness fails for a specific $d_{i}$ if the adversary controls $P_{i}$ (preventing the first condition) and controls at least kt+1k-t+1 parties in $\mathcal{G}_{i}$ (preventing the second condition) during the reconstruction phase.
Theorem 2 (Liveness of Reconstruction).
Assume the FDKG generation phase completed successfully (realizing $\mathcal{F}_{FDKG}$). Assume the NIZK proofs used in the reconstruction phase ($\pi_{PS_{i}}$, $\pi_{SPS_{j,i}}$) are computationally sound. The FDKG protocol achieves Liveness of Reconstruction if, for every participant $P_{i}\in\mathbb{D}$, the adversary does not control both $P_{i}$ and kt+1k-t+1 guardians in $\mathcal{G}_{i}$ during the reconstruction phase.
Proof Sketch.
The proof relies on the properties of SSS and the soundness of the NIZK proofs $\pi_{PS_{i}}$ and $\pi_{SPS_{j,i}}$. For any $P_{i}\in\mathbb{D}$:
  1. 1.
    If $P_{i}$ is honest/available in Round 2, it broadcasts $d_{i}$ and $\pi_{PS_{i}}$. Soundness ensures $\pi_{PS_{i}}$ only verifies if $d_{i}$ corresponds to $E_{i}$.
  2. 2.
    If $P_{i}$ is corrupt/unavailable, reconstruction depends on its guardians $\mathcal{G}_{i}$. If the adversary controls fewer than kt+1k-t+1 guardians in $\mathcal{G}_{i}$, then at least $t$ guardians are honest/available. Each such honest guardian $P_{j}\in\mathcal{G}_{i}$ will decrypt $C_{i,j}$ to get $s_{i,j}$ using $\textrm{sk}_{j}$, and broadcast $s_{i,j}$ with $\pi_{SPS_{i,j}}$. Soundness ensures $\pi_{SPS_{i,j}}$ only verifies if $s_{i,j}$ is the correct decryption of $C_{i,j}$ using $\textrm{sk}_{j}$. Correctness of Round 1 (guaranteed by soundness of $\pi_{FDKG_{i}}$) ensures $C_{i,j}$ encrypted $f_{i}(j)$. Thus, the observer obtains at least $t$ valid shares $\{s_{i,j}=f_{i}(j)\}$. By SSS properties, these $t$ shares allow reconstruction of $d_{i}=f_{i}(0)$.
The theorem’s condition ensures that at least one of these two recovery methods is available for each $d_{i}$, thus guaranteeing $\textbf{d}=\sum d_{i}$ can be computed. ∎
VI Liveness Simulations
To evaluate FDKG’s resilience under realistic deployment conditions, we conducted simulations exploring its performance under various parameter settings, comparing its liveness properties against expectations from traditional DKG under similar failure models. Our analysis focuses on three aspects: 1) parameter sensitivity, 2) scalability across network sizes, and 3) topology-dependent performance characteristics. The primary objective was to determine which configurations optimize liveness under varying conditions, thereby offering insights for practical deployment.
VI-A Simulation Methodology
We developed a simulation framework implementing both FDKG and standard DKG dynamics across two network topologies. Each simulation involves the following steps:
  1. 1.
    Network Generation: We considered two network topologies to represent different communication patterns:
    • Barabási-Albert (BA): A scale-free model designed to simulate scenarios where nodes exhibit preferential attachment, creating highly connected hubs.
    • Random Graph: A uniform model representing scenarios where nodes interact uniformly, with no highly connected hubs.
    Networks were initialized with a specified number of nodes ($n$). Guardian sets of size $k$ were then selected for participating nodes based on the specific topology rules (preferential attachment for BA, uniform for Random Graph).
  2. 2.
    Participant Selection: For each simulation run, subsets of nodes were selected to participate in the protocol phases:
    • Distribution Participants: A fraction $p$ of nodes were selected to participate, using either uniform random sampling (for the Random Graph model) or degree-weighted sampling (for the Barabási-Albert model). Let this set be $\mathbb{D}$.
    • Reconstruction Participants: A fraction $r$ of the initial Distribution Participants (from $\mathbb{D}$) were randomly selected to be available for the Reconstruction Phase, representing the retention rate. Let this set be $\mathbb{T}$.
  3. 3.
    Liveness Check: For each simulation run, the protocol checked whether the set of available reconstruction participants ($\mathbb{T}$) could successfully meet the conditions required for reconstructing the secret, respecting the FDKG rules. Success was defined as achieving Liveness of Reconstruction (Definition 1) despite node failures or absences modelled by the selection process above. Specifically, for each $P_{i}\in\mathbb{D}$, we checked if either $P_{i}\in\mathbb{T}$ or if at least $t$ guardians from $\mathcal{G}_{i}$ were present in $\mathbb{T}$. If this condition held for all $i\in\mathbb{D}$, the run was marked as successful.
The web simulator is available at: https://fdkg.stan.bar, source code is available at https://github.com/stanbar/fdkg. Figure 4 presents a snapshot of two different network topologies for the same setting.
Figure 4: Comparison of Random (left) and Barabási-Albert (right) guardian set selection policies for $n=100,p=0.3,r=0.9,k=5,t=3$. Gray nodes indicate the absence of the node, green nodes represent those that were present in both the distribution and reconstruction phases, blue nodes were present only in the first (distribution) phase, and orange nodes only in the second (reconstruction) phase.
VI-B Parameter Space and Simulation Workflow
We explored a comprehensive range of parameters to capture diverse deployment scenarios:
  • Number of Participants ($n$): $\{50,100,200,\dots,1000\}$.
  • Participation Rate ($p$): $\{0,1,0.2,\dots,0.9,1.0\}$.
  • Retention Rate ($r$): $\{0.1,0.2,\dots,1.0\}$.
  • Number of Guardians ($k$): $\{1,3,\dots,n-1\}$.
  • Threshold ($t$): Takes values from $1$ up to $k$.
For each unique combination of parameters, 100 independent simulations were executed, resulting in over 100 million total runs.
Figure 5 shows the scalability comparison between network models with fixed $p=1.0$, and $r=0.7$. The Y axis (success rate) is a mean value among all the $(k,t)$ configurations. Standard DKG corresponds to the FDKG case where $p=1.0$ and k=n1k=n-1 (all other nodes are guardians), thus its performance represents an upper bound on the mean success rate for a given $r$. In standard DKG each node connects to every other node, creating a complete graph, so there is no difference between BA and Random guardian selection policies. The BA policy (solid lines) generally outperforms the random policy (dashed lines) due to hub resilience, especially at smaller $n$.
Figure 5: Scalability comparison between average guardian set size and different policy selection with fixed $p=1.0,r=0.7$. The blue line represents the standard DKG scheme (k=n1k=n-1).
We then break down the mean values to see how each parameter influences the success rate. To focus on the impact of $p,r,k,t$, we subsequently use only the results from the BA policy selection.
Figure 6 shows success rates for each of the $(k,t)$ configurations for fixed $n=100$, $p=0.8$. Figure 6(a) uses $r=0.5$, while Figure 6(b) uses $r=0.9$, highlighting the critical significance of the retention parameter $r$.
Figure 6: Heatmaps of success rates by guardian configuration for fixed $n=100$, $p=0.8$, $r=0.5$ (top) and $r=0.9$ (bottom). Color intensity represents success probability from 0 (yellow) to 1 (dark blue).
Figure 7 reveals three critical relationships through our multivariate analysis. A higher retention rate ($r=0.9$) guarantees a high success rate even for a smaller number of guardians ($k=10$) at a partial participation rate ($p=0.75$). A lower retention rate ($r=0.5$) may require a higher participation rate ($\geq 0.6$) and a higher number of guardians ($k\geq 40$) to achieve a perfect success rate. Lower retention rates ($\leq 0.5$) combined with a low number of guardians ($k\leq 10$) may not guarantee high success rates even at perfect participation ($p=1.0$).
Figure 7: Success rate vs participation rate (top) and retention rate (bottom) for varying $k$, and $p$, or $r$ combinations, with fixed $n=100$ and $t=k/4$. Line styles represent retention or participation rates, colors indicate guardian set sizes.
VI-C Results
The results reveal clear trends in the factors influencing the success rate (liveness probability). The participation rate ($p$) has a significant impact, with higher participation rates consistently improving success rates. This is because more nodes actively contributing in Round 1 increases the likelihood that enough participants or their guardians are available for reconstruction.
Retention rate ($r$) also plays a crucial role. Higher retention rates both reduce the reliance on guardians and increase the chances of meeting the threshold during reconstruction phase. For instance, when $r=1.0$, the protocol always achieves Liveness even with very low participation $p$ and guardians $k$, it is because the party who contributed to the first phase always reconstruct the share in the second phase. However, such scenarios were excluded from the analysis since they do not represent the challenges of FDKG systems.
The number of guardians ($k$) contributes to the robustness of the system, especially in scenarios with a larger total number of nodes. Larger guardian sets provide greater redundancy, mitigating the impact of absent nodes during reconstruction. This becomes increasingly important as the system scales, where the probability of absences rises proportionally with the total number of participants.
Threshold values ($t$) directly affect liveness. Lower thresholds (requiring fewer available guardians) are naturally more likely to succeed. However, higher thresholds can still succeed if combined with sufficiently large guardian sets ($k$), high participation ($p$), and high retention ($r$).
The results highlight also the impact of guardians selection policy on liveness, particularly the advantage of the Barabási-Albert model over the Random model. The superior performance of the Barabási-Albert model is primarily due to its weighted selection policy. In this policy, nodes with higher connectivity (hubs) are more likely to be selected as participants in both the Distribution and Reconstruction phases. This approach mirrors real-world social structures where highly connected individuals or entities play a more central role in communication and collaboration. Consequently, these hubs are not only more resilient to failures, but are also more likely to engage in critical democratic processes, such as voting or key reconstruction. However, the results show that with the higher network sizes the difference between those models diminishes.
To assess FDKG’s practical performance under realistic deployment conditions—such as node unavailability, varying participation rates, and network churn—we conducted simulations. These simulations complement the formal security guarantees provided in Section V, which ensure resilience against adversarial attacks under cryptographic assumptions. Importantly, the simulations do not aim to prove security; rather, they illustrate how protocol parameters like participation rate ($p$), retention rate ($r$), guardian set size ($k$), and threshold ($t$) impact the likelihood of successful key reconstruction in scenarios with non-adversarial failures (e.g., honest node crashes or dropouts). The results offer practical insights for parameter selection in real-world deployments, where both adversarial and non-adversarial challenges must be considered.
VII Application: FDKG-Enhanced Voting Protocol
In this section, we demonstrate an application of FDKG within an internet voting protocol. This protocol builds upon the foundational three-round structure common in threshold cryptosystems [33], incorporates multi-candidate vote encoding [34], and crucially utilizes FDKG (Section IV) for decentralized and robust key management.
The core idea is to use FDKG to establish a threshold ElGamal public key E where the corresponding secret key d is distributed among participants and their guardians, mitigating single points of failure and handling dynamic participation.
The voting protocol consists of three primary rounds:
  1. 1.
    Key Generation (FDKG): Participants establish the joint public encryption key E and distribute shares of their partial decryption keys.
  2. 2.
    Vote Casting: Voters encrypt their choices under E and submit their ballots with proofs of validity.
  3. 3.
    Tallying: Participants collaboratively decrypt the sum of encrypted votes to reveal the final tally.
VII-A Round 1: Federated Distributed Key Generation
This round follows the $\Pi_{FDKG}$ protocol exactly as described in Section IV-B1.
  • Participation: Optional for any party $P_{i}\in\mathbb{P}$.
  • Actions: Each participating party $P_{i}$ selects its guardian set $\mathcal{G}_{i}$, generates its partial key pair $(d_{i},E_{i})$, computes and encrypts shares $C_{i,j}$ for its guardians, generates a NIZK proof $\pi_{FDKG_{i}}$, and broadcasts its contribution.
  • Outcome: Observers verify the proofs $\pi_{FDKG_{i}}$ to determine the set of valid participants $\mathbb{D}$. The global public encryption key for the election is computed as $\textbf{E}=\prod_{P_{i}\in\mathbb{D}}E_{i}$. The public state includes $\mathbb{D}$, E, and the broadcasted tuples $\{E_{i},\mathcal{G}_{i},\{C_{i,j}\},\pi_{FDKG_{i}}\}_{P_{i}\in\mathbb{D}}$.
VII-B Round 2: Vote Casting
Let $\mathbb{V}\subseteq\mathbb{P}$ be the set of eligible voters participating in this round. For each voter $P_{i}\in\mathbb{V}$:
  1. 1.
    Vote Encoding: Encode the chosen candidate into a scalar value $v_{i}\in\mathbb{Z}_{q}$. We use the power-of-2 encoding from [35]:
    $v_{i}\ =\ \begin{cases}G^{2^{0}}&\text{for candidate 1}\\ G^{2^{m}}&\text{for candidate 2}\\ \vdots&\vdots\\ G^{2^{(c-1)m}}&\text{for candidate }c\end{cases}$
    where $m$ is the smallest integer s.t. $2^{m}>n$, and $c$ is the number of candidates. The vote is represented as a point on the curve.
  2. 2.
    Vote Encryption (ElGamal): Generate a random blinding factor $r_{i}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}$. Compute the encrypted ballot $B_{i}=(C1_{i},C2_{i})$ where:
    $C1_{i}$ $=G^{r_{i}}$
    $C2_{i}$ $=\textbf{E}^{r_{i}}\cdot v_{i}\quad\text{(using group operation in }G)$
  3. 3.
    Proof Generation: Generate a NIZK proof $\pi_{Ballot_{i}}$ demonstrating that $B_{i}$ is a valid ElGamal encryption of a vote $v_{i}$ from the allowed set $\{G^{2^{0}},G^{2^{m}},\dots,G^{2^{(c-1)m}}\}$ under public key E, without revealing $v_{i}$ or $r_{i}$. (See Appendix A-D).
  4. 4.
    Broadcast: Broadcast the ballot and proof $(B_{i},\pi_{Ballot_{i}})$.
State after Voting
The broadcast channel is appended with the set of valid ballots and proofs:
  • $\{(B_{i},\pi_{Ballot_{i}})\mid P_{i}\in\mathbb{V}\text{ and }\pi_{Ballot_{i}}\text{ is valid}\}$. Let $\mathbb{V}^{\prime}\subseteq\mathbb{V}$ be the set of voters whose ballots verified.
VII-C Round 3: Tallying
This round involves threshold decryption of the combined ballots. It follows the structure of the FDKG Reconstruction Phase (Section IV-B2).
VII-C1 Online Tally
A subset of parties $\mathbb{T}\subseteq\mathbb{P}$ participate. For each party $P_{i}\in\mathbb{T}$:
  1. 1.
    Compute Homomorphic Sum: Aggregate the first components of all valid ballots: $C1=\prod_{P_{j}\in\mathbb{V}^{\prime}}C1_{j}$.
  2. 2.
    Provide Partial Decryption (if applicable): If $P_{i}\in\mathbb{D}$:
    • Compute the partial decryption: $\mathrm{PD}_{i}=(C1)^{d_{i}}$.
    • Generate a NIZK proof $\pi_{PD_{i}}$ proving that $\mathrm{PD}_{i}$ was computed correctly using the secret $d_{i}$ corresponding to the public $E_{i}$. (See Appendix A-E).
    • Broadcast $(P_{i},\text{pdecrypt},\mathrm{PD}_{i},\pi_{PD_{i}})$.
  3. 3.
    Provide Decrypted Shares: For each encrypted share $C_{j,i}$ received in Round 1 (where $P_{j}\in\mathbb{D}$ and $P_{i}\in\mathcal{G}_{j}$):
    • Decrypt the share: $s_{j,i}=\texttt{Dec}_{\textrm{sk}_{i}}(C_{j,i})$.
    • Compute the share’s contribution to decryption: $\mathrm{PD}_{j}]_{i}=(C1)^{s_{j,i}}$.
    • Generate a NIZK proof $\pi_{PDS_{j,i}}$ proving correct decryption of $C_{j,i}$ to $s_{j,i}$ and correct exponentiation of $C1$ by $s_{j,i}$. (See Appendix A-F).
    • Broadcast $(P_{i},\text{pdecryptshare},P_{j},[\mathrm{PD}_{j}]_{i},\pi_{PDS_{j,i}})$.
State after Online Tally
The broadcast channel contains potentially revealed partial decryptions $(\mathrm{PD}_{i},\pi_{PD_{i}})$ and shares of partial decryptions $([\mathrm{PD}_{j}]_{i},\pi_{PDS_{j,i}})$ from parties in $\mathbb{T}$.
VII-C2 Offline Tally
Performable by any observer using publicly available data.
  1. 1.
    Aggregate Ballot Components: Compute the aggregated components of valid ballots:
    $C1$ $=\prod_{P_{k}\in\mathbb{V}^{\prime}}C1_{k}$
    $C2$ $=\prod_{P_{k}\in\mathbb{V}^{\prime}}C2_{k}$
  2. 2.
    Reconstruct Partial Decryptions: For each $P_{i}\in\mathbb{D}$:
    • Try to find a valid broadcast $(P_{i},\text{pdecrypt},\mathrm{PD}_{i},\pi_{PD_{i}})$. If found and $\pi_{PD_{i}}$ verifies, use $\mathrm{PD}_{i}=\mathrm{PD}_{i}$.
    • Otherwise, collect all valid broadcasted shares $(P_{k},\text{pdecryptshare},P_{i},[\mathrm{PD}_{i}]_{j},\pi_{PDS_{i,j}})$ where $P_{k}=P_{j}$ and $\pi_{PDS_{i,j}}$ verifies. Let $I_{i}$ be the set of indices $j$ for which valid shares were found.
    • If $|I_{i}|\geq t$: Select $I^{\prime}_{i}\subseteq I_{i}$, $|I^{\prime}_{i}|=t$. Use Lagrange interpolation in the exponent to compute $\mathrm{PD}_{i}=\prod_{P_{j}\in I^{\prime}_{i}}([\mathrm{PD}_{i}]_{j})^{\lambda_{j}}$. Verify this reconstructed value against $E_{i}$.
    • If neither direct secret nor enough shares are available/valid, the tally fails.
  3. 3.
    Combine Partial Decryptions: Compute the combined decryption factor $Z$:
    $Z=\prod_{P_{i}\in\mathbb{D}}\mathrm{PD}_{i}=(C1)^{\sum d_{i}}=(C1)^{\textbf{d}}$
  4. 4.
    Decrypt Final Result: Compute the product of the encoded votes:
    $M=C2\cdot Z^{-1}=\left(\prod_{P_{k}\in\mathbb{V}^{\prime}}\textbf{E}^{r_{i}}\cdot v_{k}\right)\cdot\left(\prod_{P_{k}\in\mathbb{V}^{\prime}}G^{r_{i}}\right)^{-\textbf{d}}$
    Since $\textbf{E}=G^{\textbf{d}}$, this simplifies to:
    $M=\prod_{P_{k}\in\mathbb{V}^{\prime}}v_{k}=G^{\sum x_{k}2^{(k-1)m}}$
    where $x_{k}$ is the number of votes for candidate $k$.
  5. 5.
    Extract Counts: Solve the discrete logarithm problem $M=G^{\text{result}}$. Since ‘result‘ = $\sum x_{k}2^{(k-1)m}$ and the total number of votes $|\mathbb{V}^{\prime}|$ is known (bounding the $x_{k}$), the counts $x_{k}$ can be extracted efficiently, e.g., using Pollard’s Rho or Baby-Step Giant-Step for the small resulting exponent, or simply by examining the binary representation if $m$ is chosen appropriately [34].
VIII Performance Evaluation
This section presents a performance analysis of the FDKG-enhanced voting protocol described in Section VII. We evaluated computational costs (NIZK proof generation times and offline tally time) and communication costs (total broadcast message sizes). All NIZK evaluations use Groth16 [23] as the NIZK proof system, assuming a proof size of 256 bytes (uncompressed). We briefly investigated using the PLONK [24] proof system as an alternative potentially avoiding a trusted setup, but found it computationally infeasible for our circuits; for instance, generating a $\pi_{FDKG_{i}}$ proof with parameters $(t=3,k=10)$ took over 8 minutes with PLONK, and larger configurations exceeded available memory resources. Therefore, we proceeded with Groth16.
VIII-A Computational and Communication Costs per Message Type
Table II summarizes the core performance metrics for the different messages broadcast during the online phases of the protocol. It includes the time to generate the required NIZK proof and the total message size, which comprises the cryptographic payload plus the 256-byte Groth16 proof.
Table II: Performance Summary: NIZK Proving Time and Total Message Size per Broadcast Message Type.
Measurement FDKG Distribution Ballot Broadcast Partial Dec. Broadcast Part. Dec. Share Broadcast
(t=3, k=10) (t=10, k=30) (t=30, k=100)
NIZK Proving Time 3.358 s 4.415 s 14.786 s 0.211 s 0.619 s 0.580 s/share
Base Payload Size 1664 B 4864 B 16064 B 128 B 64 B 64 B/share
NIZK Proof Size 256 B 256 B 256 B 256 B 256 B 256 B/share
Total Message Size (Bytes) 1920 5120 16320 384 320 320/share
The online computational cost is dominated by the generation of the $\pi_{FDKG_{i}}$ proof, especially as $t$ and $k$ increase. The communication cost for the FDKG distribution phase is primarily driven by $k$. Messages related to voting and tallying are significantly smaller individually, but the total communication cost of tallying depends heavily on how many secrets and shares need to be revealed publicly.
VIII-B Offline Tally: Vote Extraction Cost
The final step of the Offline Tally phase (Section VII, Offline Tally, Step 5) involves extracting the individual vote counts $\{x_{k}\}$ from the aggregated result $M=G^{\sum x_{k}2^{(k-1)m}}$. This requires solving a Discrete Logarithm Problem (DLP) where the exponent has a specific known structure. Since the maximum value of the exponent is bounded by the total number of valid votes $|\mathbb{V}|$, standard algorithms like Baby-Step Giant-Step or Pollard’s Rho can be used.
We measured the time required for this vote extraction step for varying numbers of voters $|\mathbb{V}|$ and candidates $c$. The results are presented in Figure 8. The data shows that the time required grows approximately linearly with the number of voters (for a fixed number of candidates) but grows exponentially with the number of candidates $c$. For instance, with 2 candidates, extracting results for up to 1000 voters takes only milliseconds. However, with 5 candidates, extracting the result for 100 voters took over 100 seconds.
This indicates that while the FDKG and online tallying phases scale well, the vote extraction step using this specific power-of-2 encoding [35] imposes a practical limit on the number of candidates that can be efficiently supported, especially when combined with a large number of voters.
Figure 8: Time required to solve DLP for vote extraction with respect to the number of candidates and number of voters (log-log scale).
VIII-C Example Scenario Calculation
Let’s estimate the total broadcast communication size for a specific scenario:
  • $n=100$ potential participants.
  • $|\mathbb{D}|=50$ valid participants in Round 1 (FDKG Distribution).
  • Each participant uses $k=40$ guardians.
  • $|\mathbb{V}|=50$ valid ballots cast in Round 2.
  • In Round 3 (Online Tally), assume $|\mathbb{T}\cap\mathbb{D}|=40$ participants reveal their secrets directly.
  • Assume for reconstruction, each of the 40 online talliers reveals, on average, 40 shares, this leads to approximately $40\times 40=1600$ total shares being revealed (this is a simplification, the actual number depends on guardian set overlaps and the specific set of online talliers).
Using the ”Total Message Size (Bytes)” figures from Table II:
  1. 1.
    FDKG Dist.: $50\times(64+40\times 160+256)=50\times 6720\text{ B}=336000$ B (336.0 kB).
  2. 2.
    Voting: $50\times 384\text{ B}=19200$ B (19.2 kB).
  3. 3.
    Tallying (Online):
    • Secrets: $40\times 320\text{ B}=12800$ B (12.8 kB).
    • Shares: $1600\times 320\text{ B}=512000$ B (512.0 kB).
  4. 4.
    Total Broadcast: $336000+19200+12800+512000=880000$ B (880.0 kB).
The offline tally time for this scenario (50 voters) would be milliseconds for 2-4 candidates, but could increase to several minutes for 5 or more candidates, as indicated by Figure 8.
IX Deployments
The FDKG protocol (and FDKG-based voting application), requiring only a shared communication space (message board) where participants can publicly post and read messages, is adaptable to various deployment environments. We consider three illustrative scenarios:
  1. 1.
    Public Blockchains: The protocol can be implemented on public blockchains. For instance, interactions could be managed via smart contracts. Features like ERC-4337 (Account Abstraction) [36] could further simplify deployment by enabling centralized payment of transaction fees, potentially lowering adoption barriers for participants.
  2. 2.
    Peer-to-peer Networks: The protocol is inherently suitable for peer-to-peer (P2P) networks. Technologies such as Wesh Network [^2] Wesh Network, an asynchronous mesh network protocol by Berty Technologies, https://wesh.network/, which provide decentralized communication infrastructure, can serve as a foundation for FDKG deployments.
  3. 3.
    Centralized Messengers: For ease of deployment and accessibility, FDKG could be layered on top of existing centralized messenger platforms (e.g., Telegram, Signal, WhatsApp), where a group chat acts as the message board. However, it is important to note that the centralized nature of the messenger could introduce a trusted third party capable of censorship, potentially impacting protocol liveness or fairness if it selectively blocks messages.
Figure 9 visually depicts these potential deployment architectures.
Figure 9: Three possible deployments of the protocol: Centralised messenger, ad-hoc peer-to-peer network, and public blockchain.
X Limitations and Future Work
While FDKG offers enhanced flexibility and resilience to dynamic participation, our analysis and implementation highlight areas for further improvement and research.
Cryptographic Overhead: The protocol inherits computational overhead from its cryptographic building blocks, particularly NIZK proofs. Although Groth16 proofs, as used in our implementation, are efficient for mid-scale settings (e.g., tens to a few hundred participants), very large-scale or time-critical applications (like high-throughput elections) might necessitate faster proving systems (e.g., different SNARKs or STARKs) or techniques for proof aggregation to reduce verification load.
Offline Tallying Scalability: As discussed in Section VIII, the offline tallying step in our FDKG-enhanced voting example, specifically the brute-force decoding of the final aggregated vote via DLP solution, exhibits exponential scaling with the number of candidates. Developing more efficient multi-candidate vote encoding schemes or advanced number-theoretic techniques for vote extraction remains an important open research challenge for such applications.
Guardian Set Selection: Our liveness simulations (Section VI) explored heuristics like Barabási-Albert for guardian selection, which showed promise. However, a deeper investigation into context-aware guardian selection strategies could further bolster FDKG’s practical reliability. This might involve incorporating social trust metrics, node reputation systems, or application-specific criteria to optimize for liveness and security against targeted attacks.
Trusted Setup for NIZKs: Our current NIZK instantiation (Groth16) requires a trusted setup for its Common Reference String (CRS). While this is a one-time cost per circuit, it represents a potential point of centralization or trust. Future work could explore integrating FDKG with NIZK systems that have transparent or universal setups (e.g., STARKs, or SNARKs like PLONK with a universal updatable SRS) to mitigate this dependency, although this may involve performance trade-offs.
Broader Applications: Although this paper focuses on an i-voting application, the FDKG concept is a general cryptographic primitive. Its applicability extends to any setting requiring robust, threshold-based secret reconstruction where global membership is uncertain or dynamic. Potential areas include collaborative data encryption/decryption in ad-hoc networks, decentralized credential management systems, and key management for emergent governance structures like Decentralized Autonomous Organizations (DAOs). Exploring these diverse applications and tailoring FDKG to their specific needs presents a rich avenue for future work.
Overall, these points highlight opportunities for optimizing cryptographic components, enhancing theoretical models (e.g., for adaptive adversaries in the FDKG context), and broadening the practical deployment of FDKG.
XI Discussion and Conclusions
This paper introduced Federated Distributed Key Generation (FDKG), a novel protocol that enhances the flexibility and resilience of distributed key establishment in dynamic environments. Unlike standard DKG protocols that assume a fixed set of participants and a global threshold, FDKG empowers individual participants to select their own ”guardian sets” of size $k$ for share recovery, accommodating scenarios with unknown or fluctuating network membership. Our security analysis (Section V) demonstrates that FDKG achieves correctness and privacy for the key generation phase, and liveness for key reconstruction, provided an adversary does not control both a participant and kt+1k-t+1 of its chosen guardians during reconstruction.
FDKG generalizes traditional DKG by integrating principles of federated trust, reminiscent of Federated Byzantine Agreement (FBA) systems like the Stellar Consensus Protocol. This allows for dynamic trust delegation at the participant level. By achieving its core functionality in two communication rounds (distribution and reconstruction), FDKG is suitable for systems involving human interaction, such as the i-voting application presented. The reliance on NIZK proofs (e.g., zk-SNARKs) for verifiability ensures protocol integrity but introduces computational overhead and, in the case of Groth16, a trusted setup requirement, which are important considerations for deployment. Future work could explore NIZKs with transparent setups (e.g., STARKs or Bulletproofs, though the latter might impact succinctness) to alleviate this.
Our simulations (Section VI) provided practical insights into FDKG’s operational robustness. Liveness was shown to be heavily influenced by participant engagement (participation rate $p$) and continued availability (retention rate $r$). For example, in a network of $n=100$ parties with $p=0.8$ and $r=0.9$, FDKG consistently achieved liveness for thresholds $t<0.7k$. However, with a lower retention of $r=0.5$, perfect liveness was only achieved for smaller thresholds $t\leq 0.4k$ (Figure 6). The choice of guardian selection topology also matters, with Barabási-Albert (BA) providing a slight resilience advantage over random selection in smaller networks due to its hub-centric nature (Figure 5), although this advantage diminishes for $n>1000$.
Performance evaluations (Section VIII) quantified the computational and communication costs. For $n=100$ participants with $k=40$ guardians, the FDKG distribution phase incurred approximately 336 kB of broadcast data per participant. NIZK proof generation, while enabling verifiability, is the main computational bottleneck, with $\pi_{FDKG_{i}}$ proving times ranging from $3.36$s (for $k=10$) to $14.79$s (for $k=100$) in our experiments (Table II). These figures highlight the trade-off between the flexibility offered by larger guardian sets and the associated costs, underscoring the need for careful parameter tuning based on specific application constraints: smaller $(k,t)$ values may be preferable for latency-sensitive systems, while larger values enhance resilience in more adversarial settings. A critical aspect of FDKG is balancing guardian set size ($k$) against security; while smaller sets reduce overhead, they concentrate trust, making individual guardians more critical for liveness if the original participant is unavailable or malicious.
We envision FDKG as a versatile cryptographic primitive for multi-party systems that require robust and verifiable key distribution but must operate under conditions of incomplete or dynamic participation. Future research directions include optimizing NIZK proofs for FDKG-specific relations, refining vote decoding mechanisms for FDKG-based voting schemes, and developing more sophisticated, context-aware guardian selection strategies to further enhance FDKG’s practical usability and security. By embracing a decentralized and adaptive ethos, FDKG offers a pathway to building scalable and secure distributed systems for diverse collaborative environments.
Reproducible Research
To ensure the reproducibility of our findings, all source code, datasets, and scripts used to generate results are publicly available. The repository is hosted at https://github.com/stanbar/fdkg.
Acknowledgment
We thank Lev Soukhanov for his early contributions to the FDKG concept and mechanics, and Maya Dotan, Jonathan Katz, Oskar Goldhahn, and Thomas Haines for their insightful reviews, enhancing the paper’s structure and cryptographic rigor.
Appendix A zkSNARK Constructions
This section details the specific statements and witnesses for the NIZK proofs used in the FDKG protocol and the associated voting application. We assume a NIZK proof system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ operating with a common reference string crs over a suitable group $\mathbb{G}$ with generator $G$ and order $q$. The proofs are instantiated using Groth16 [23]. Each subsection defines the relation $\mathcal{R}$ implicitly through the statement $x$ and witness $w$. Proofs are generated as $\pi\leftarrow\textrm{Prove}(\textrm{crs},x,w)$ and verified by checking $\textrm{Verify}(\textrm{crs},x,\pi)\stackrel{{\scriptstyle?}}{{=}}1$.
The detailed Circom circuit implementations for these proofs are publicly available on https://github.com/stanbar/FDKG.
A-A FDKG Proof ($\pi_{FDKG_{i}}$)
Proves correct generation and encryption of shares by participant $P_{i}$ during the distribution phase.
  • Statement ($x_{i}$): Consists of the party’s partial public key $E_{i}$, the set of PKE public keys $\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}}$ for its guardians, and the set of corresponding encrypted shares $\mathbb{C}_{i}=\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}}$.
  • Witness ($w_{i}$): Consists of the coefficients $\{a_{0},\dots,a_{t-1}\}$ of $P_{i}$’s secret polynomial $f_{i}(X)$ (where $d_{i}=a_{0}=f_{i}(0)$), and the PKE randomness $\{(k_{i,j},r_{i,j})\}_{P_{j}\in\mathcal{G}_{i}}$ used to produce each ciphertext in $\mathbb{C}_{i}$.
  • Relation $\mathcal{R}_{FDKG}$ ($((x_{i}),w_{i})\in\mathcal{R}_{FDKG}$): The circuit (PVSS.circom) checks that:
    1. 1.
      The public $E_{i}$ is correctly derived from the witness coefficient $a_{0}$ (i.e., $E_{i}=G^{a_{0}}$).
    2. 2.
      For each guardian $P_{j}\in\mathcal{G}_{i}$:
      • The plaintext share $s_{i,j}=f_{i}(j)$ is correctly evaluated from the witness polynomial coefficients $\{a_{k}\}$.
      • The ciphertext $C_{i,j}\in\mathbb{C}_{i}$ is a valid ElGamal variant encryption of the plaintext share $s_{i,j}$ under guardian $P_{j}$’s PKE public key $\textrm{pk}_{j}$, using the witness randomness $(k_{i,j},r_{i,j})$. This involves verifying the structure $C_{1}=G^{k_{i,j}}$, $M=G^{r_{i,j}}$, $C_{2}=(\textrm{pk}_{j})^{k_{i,j}}\cdot M$, and $\Delta=M.x-s_{i,j}$.
A-B Partial Secret Proof ($\pi_{PS_{i}}$)
Proves knowledge of the partial secret key $d_{i}$ corresponding to the partial public key $E_{i}$ during reconstruction.
A-C Share Decryption Proof ($\pi_{SPS_{j,i}}$)
Proves correct decryption of an encrypted share $C_{j,i}$ by the recipient party $P_{i}$ (guardian) using its PKE secret key $\textrm{sk}_{i}$.
  • Statement ($x_{j,i}$): Consists of the encrypted share $C_{j,i}=(C_{1},C_{2},\Delta)$, the recipient’s (guardian $P_{i}$’s) PKE public key $\textrm{pk}_{i}$, and the claimed plaintext share $s_{j,i}$.
  • Witness ($w_{j,i}$): The recipient’s (guardian $P_{i}$’s) PKE secret key $\textrm{sk}_{i}$.
  • Relation $\mathcal{R}_{SPS}$ ($((x_{j,i}),w_{j,i})\in\mathcal{R}_{SPS}$): The circuit (decrypt_share.circom) checks that applying the ElGamal variant decryption steps using $\textrm{sk}_{i}$ to $(C_{1},C_{2},\Delta)$ yields the plaintext $s_{j,i}$. That is, it computes $M=C_{2}\cdot(C_{1}^{\textrm{sk}_{i}})^{-1}$ and verifies $M.x-\Delta=s_{j,i}\pmod{q}$.
A-D Ballot Proof ($\pi_{Ballot_{i}}$)
Proves a ballot $B_{i}$ is a valid ElGamal encryption of an allowed vote $v_{i}$ under the global public key E.
  • Statement ($x_{B_{i}}$): The global public key E and the ballot $B_{i}=(C1_{i},C2_{i})$.
  • Witness ($w_{B_{i}}$): The blinding factor $r_{i}\in\mathbb{Z}_{q}$ and the encoded vote message $M_{vote}$.
  • Relation $\mathcal{R}_{Ballot}$ ($((x_{B_{i}}),w_{B_{i}})\in\mathcal{R}_{Ballot}$): The circuit (encrypt_ballot.circom) checks that:
    1. 1.
      $C1_{i}=G^{r_{i}}$.
    2. 2.
      $C2_{i}=\textbf{E}^{r_{i}}\cdot G^{M_{vote}}$.
    3. 3.
      The witness $M_{vote}$ corresponds to one of the allowed vote encodings (e.g., $M_{vote}\in\{2^{0},\dots,2^{(c-1)m}\}$).
A-E Partial Decryption Proof ($\pi_{PD_{i}}$)
Proves correct computation of the partial decryption $\mathrm{PD}_{i}$ by party $P_{i}$.
  • Statement ($x_{PD_{i}}$): The aggregated ElGamal component $C1$, the computed partial decryption $\mathrm{PD}_{i}$, and the party’s partial public key $E_{i}$.
  • Witness ($w_{PD_{i}}$): The party’s partial secret key $d_{i}$.
  • Relation $\mathcal{R}_{PD}$ ($((x_{PD_{i}}),w_{PD_{i}})\in\mathcal{R}_{PD}$): The circuit (partial_decryption.circom) checks that:
    1. 1.
      $E_{i}=G^{d_{i}}$.
    2. 2.
      $\mathrm{PD}_{i}=(C1)^{d_{i}}$.
A-F Partial Decryption Share Proof ($\pi_{PDS_{j,i}}$)
Proves correct decryption of an SSS share $s_{j,i}$ (originally from $P_{j}$, decrypted by guardian $P_{i}$) and its application to computing $P_{i}$’s share of the partial decryption of $C1$.
  • Statement ($x_{PDS_{j,i}}$): The aggregated ElGamal component $C1$, the computed share of partial decryption $\mathrm{PD}_{j}]_{i}$, the encrypted SSS share $C_{j,i}=(C_{1},C_{2},\Delta)$, and the decrypting guardian $P_{i}$’s PKE public key $\textrm{pk}_{i}$.
  • Witness ($w_{PDS_{j,i}}$): The decrypting guardian $P_{i}$’s PKE secret key $\textrm{sk}_{i}$.
  • Relation $\mathcal{R}_{PDS}$ ($((x_{PDS_{j,i}}),w_{PDS_{j,i}})\in\mathcal{R}_{PDS}$): The circuit (partial_decryption_share.circom) first internally computes the plaintext SSS share $s_{j,i}=\texttt{Dec}_{\textrm{sk}_{i}}(C_{j,i})$ (verifying the decryption step) and then checks that $\mathrm{PD}_{j}]_{i}=(C1)^{s_{j,i}}$.